If the activation energy for a given compound is found to be 22.0 kJ/mol, with a frequency factor of 8.0 × 1010 s-1, what is the rate constant for this reaction at 298 K?
Answer #1

Given:

T= 298 K

A = 8.0*10^10 s-1

Ea = 22.0 KJ/mol

= 22000 J/mol

R = 8.314 J/mol.K

use:

K = A*e^(-Ea/RT)

= 8.0*10^10*e^(-22000/(8.314*298))

= 8.*10^10*e^(-8.8797)

= 8.0*10^10*1.392*10^-4

= 1.1*10^7 s-1

Answer: 1.1*10^7 s-1